find an equation of the line that passes through (0,2)and is parallel to y=1/4x+6

Answer by meetkisai(17)

# standard line equation y = mx + c where m is slope and c is the y-intercept
since the line is parallel to y = 1/4x + 6
=> m = 1/4
and since line passes through (0, 2)
2 = 1/4*0 + c
=> c = 2
hence equation of the line is
y = 1/4x + 2